# Improving the Reasoning of Multi-Image Grounding in MLLMs via Reinforcement Learning

### v1

@@ 4 Experiments
Models Spontaneous Grounding Referential Grounding AVG
Difference Similarity Visual Reference Textual Visual+Textual
Static Robust Common OT MV Region Refer GG Reason Co-Re
Human Performance
Human 99.50* 97.87 98.00* 100.00 96.88 100.00* 98.99 91.06* 92.08 97.44 97.18
70B-Scale MLLMs
LLaVA-OV-72B 13.26 5.34 26.84 12.91 7.64 2.14 17.83 21.60 11.88 8.55 13.65
InternVL2-76B 15.91 10.64 36.40 30.73 20.83 5.74 46.46 41.28 32.67 26.50 26.72
InternVL3-78B 10.04 9.57 24.12 27.08 14.58 10.44 50.51 38.08 45.54 17.09 24.71
Qwen2-VL-72B 46.12 46.81 64.46 26.73 22.57 18.62 33.33 62.53 50.50 17.09 38.88
Qwen2.5-VL-72B 43.75 46.81 69.98 34.32 29.17 8.31 62.63 59.92 66.34 41.03 46.23
7B-Scale MLLMs
Mantis 1.52 0.00 3.31 12.18 2.08 1.00 1.01 10.02 0.00 0.85 3.20
LLaVA-OV-7B 6.06 3.19 3.43 0.18 1.04 1.08 9.09 15.43 6.93 0.85 4.73
Minicpm2.6 14.58 2.13 14.34 9.82 6.25 1.75 11.11 10.02 2.97 2.56 7.55
mPLUG-Owl3 18.56 6.38 34.93 8.55 7.64 2.41 7.07 22.85 9.09 5.98 12.35
InternVL2-8B 6.92 7.45 25.49 20.73 9.72 3.49 28.28 30.26 17.82 9.40 15.96
InternVL3-8B 23.67 14.89 47.99 14.84 6.94 12.13 7.07 34.87 16.83 2.56 18.18
Qwen2-VL-7B 27.84 38.30 19.36 20.73 11.81 25.95 23.23 58.52 48.51 11.97 28.62
Qwen2.5-VL-7B 28.03 5.32 21.81 15.23 5.56 4.07 13.13 32.26 3.96 2.56 13.19
+ SFT 53.22 44.68 82.11 26.59 35.07 39.32 79.80 59.72 53.47 23.93 49.79
+ CoT-SFT 47.35 45.74 76.59 46.14 34.03 28.93 80.81 63.93 66.34 34.19 52.41
Ours-GRPO 51.52 43.62 82.48 63.41 38.54 51.45 82.83 69.54 67.33 37.61 58.83
@@
Table 3: In-domain vs Out-of-domain performance comparison on MIG-Bench li2025migician . The in-domain subsets include static difference, common object recognition, region localization, and referential grounding, while the out-of-domain subsets comprise robust difference, object tracking, multi-view grounding, group grounding, and correspondence.
Model Single image Grounding Multi-image Grounding AVG
Ref-L4(val) Ref-L4(test) LISA(val) LISA(test) ReVOS
Qwen2.5-VL-72B 74.85 76.71 53.57 46.85 48.78 60.15
Qwen2.5-VL-7B 58.57 62.76 56.63 53.54 30.07 52.31
+ SFT 68.92 69.14 55.10 54.18 46.33 58.73
Ours 74.78 75.74 62.76 56.11 49.16 63.71
Table 4: Zero-shot evaluation on several reasoning grounding benchmarks: Ref-L4 chen2025revisiting , LISA-Grounding lai2024lisa , and ReVOS yan2024visa . The best and second-best results are marked in bold and underline, respectively.
4.1 Implementation Details
Benchmarks
We evaluate our method on the multi-image grounding benchmark MIG-Bench li2025migician , object detection dataset ODINW li2022grounded , single-image grounding datasets: Ref-L4 chen2025revisiting , LISA-Grounding lai2024lisa , and the video grounding dataset ReVOS Grounding yan2024visa . ODINW comprises 13 distinct real-world settings featuring rare object categories, assessing the model’s object perception and inference ability in practical scenarios. Moreover, to further evaluate out-of-domain generalization ability of our model, we individually employ five non-overlapping subsets from ODINW: Oxford Pets, Pascal VOC, Wildfire Smoke, Open Poetry Vision, and Vector. Ref-L4 is a large-scale referring expression benchmark with diverse object categories, instance scales, and long textual descriptions, making it ideal for assessing reasoning capabilities of MLLMs. For ReVOS, we uniformly sample six frames and task the model with grounding one frame, deriving ground-truth bounding boxes from segmentation masks originally annotated for reasoning tasks.
Training Configurations
Method OxfordPets PascalVOC WildfireSmoke OpenPoetryVision Vector AVG
@@ 4.1 Implementation Details
Table 5: Out-of-domain object detection evaluation on ODNIW li2022grounded dataset. We individually employ five non-overlapping subsets: Oxford Pets, Pascal VOC, Wildfire Smoke, Open Poetry Vision, and Vector. The best result is indicated in bold.
Closed-source Model
GPT4o 49.17 79.55 31.33 55.33 59.40 72.59 75.00 60.34
Open-source Model
Qwen2.5-VL-7B 67.50 45.45 27.33 63.33 51.88 82.96 68.02 58.07
+ SFT-63k 63.33 32.58 26.67 72.67 48.12 90.37 62.21 56.56
Ours-new 70.00 47.73 32.00 74.67 49.62 84.44 69.77 61.18
@@
We select several open-source MLLMs for comparison, such as Qwen2-VL wang2024qwen2 , Qwen2.5-VL bai2025qwen2 , InternVL2 chen2024internvl , InternVL3 zhu2025internvl3 , LLaVA-OneVision li2024llava , MiniCPM2.6 yao2024minicpm , and mPLUG-Owl3 ye2024mplug . To further verify the effectiveness of the RL paradigm, we compare our method with the SFT baseline, which is trained on the same dataset as our stage-1 CoT-SFT with a learning rate of 2e-6 and an accumulated batch size of 64.
@@
4.2 Comparison Results
MIG-Bench Comparison
As shown in Table 2, we present the comparison results in the multi-image grounding benchmark MIG-Bench li2025migician . Our method achieves state-of-the-art performance overall. Specifically, it obtains the best results across all seven subsets of referential grounding. Compared to the base model (Qwen2.5-VL-7B), our approach achieves a 45.64% improvement, and it outperforms the second-best model (Qwen2.5-VL-72B) by 12.60%, while using significantly fewer parameters.
Moreover, we conduct a comprehensive comparison of various post-training strategies on both in-domain and out-of-domain tasks, shown in Table 3. The results reveal that conventional SFT training significantly improves in-domain performance but provides limited benefits for out-of-domain generalization. CoT-SFT training strategy enhances out-of-domain generalization to a certain extent but slightly compromises in-domain accuracy. Conversely, the RL training strategy achieves superior performance, obtaining the highest scores in both in-domain and out-of-domain settings, with a notable 53.34% on out-of-domain tasks. The results demonstrate the strong generalization capability of our approach across both conventional and novel task domains.
Model High-level Obj. High-level Sub. Low-level AVG
VGR STS VCor SC FC ER CR MIC VQA FD
Closed-source Model
GPT4o 61.3 84.0 72.6 41.5 42.6 46.5 67.8 93.5 79.2 88.8 67.8
Open-source Model
Mantis 65.2 65.5 30.9 28.1 24.3 31.8 61.5 71.0 59.8 57.0 49.5
Llava-interleave 49.8 54.5 25.1 26.4 23.2 24.8 29.5 57.0 48.8 26.3 36.5
Qwen2.5-VL-7B 68.0 74.0 44.6 33.6 25.1 31.0 57.0 80.0 82.8 72.0 56.8
+ SFT 67.0 80.0 40.3 28.6 26.1 32.5 69.2 87.5 81.2 65.1 57.7
Ours 66.8 72.0 51.1 30.4 27.0 39.0 69.5 87.0 82.2 69.3 59.4
@@ 4.2 Comparison Results
Table 7: Performance on Multi-image Understanding Benchmark MMIU meng2024mmiu . High-level Obj. and High-level Sub. are represented as subtypes of high-level semantic relationship (objective) and high-level semantic relationship (subjective). The term "Low-level" denotes the subtype of low-level semantic relations. Some results are borrowed from the original paper. The best and second-best results among open-source models are marked in bold and underline, respectively.
Model Two-dimensional spatial relationship AVG
RPM JPS ICSC IQASC ITRSC ISTE HE PT SOT
Closed-source Model
GPT4o 12.5 23.5 58.5 88.0 17.5 35.0 29.5 55.0 56.0 41.7
Open-source Model
Mantis 13.5 27.5 48.0 66.0 13.0 23.5 23.0 76.0 50.0 37.8
Llava-interleave 15.0 17.5 47.5 42.0 14.0 22.5 21.5 79.0 60.5 35.5
Qwen2.5-VL-7B 13.0 26.5 55.5 75.5 7.0 29.0 23.5 58.5 55.5 37.8
+ SFT 12.5 30.5 54.0 73.0 19.5 33.5 29.5 43.0 52.5 38.6
Ours 15.0 24.5 68.0 83.5 19.0 32.5 22.0 50.0 49.5 40.4
Table 8: Performance on Multi-image Understanding Benchmark MMIU meng2024mmiu for two-dimensional spatial relationship subtype. Some results are borrowed from the original paper. The best and second-best results among open-source models are marked in bold and underline, respectively.
Zero-shot Evaluation
To further highlight the robust generalization capability of our method, we conduct zero-shot evaluation on multiple datasets spanning different tasks: the single-image grounding dataset Ref-L4 chen2025revisiting , LISA lai2024lisa , and the video grounding dataset ReVOS yan2024visa . As illustrated in Table 4, our method consistently outperforms all baselines across multiple benchmarks. Although the SFT training strategy achieves slight overall improvements, it exhibits a severe performance drop on the LISA benchmark. It sugguests that the SFT training method lacks strong generalization, relying on memorizing answers instead of learning underlying patterns. On the other hand, our RL training strategy, which is initialized with a small amount of data, achieves a substantial improvement of 4.98% over the SFT baseline and even surpasses the performance of Qwen2.5-VL-72B by 3.56%.
Out-of-domain Object Detection
The evaluation results on out-of-domain object detection task are presented in Table 5. As described in Section 4.1, we evaluate on five non-overlapping datasets from ODNIW, comprising 6,314 samples. Our full model achieves the highest average performance (54.15%) across all five datasets, surpassing the base model Qwen2.5-VL-7B by a significant improvement of 7.24%. Notably, it also outperforms both the SFT baseline and the largest scale base model, Qwen2.5-VL-72B. It demonstrate the robustness and strong generalization capability of our approach in handling out-of-domain object detection tasks.
Multi-image Understanding Evaluation
For comprehensive evaluation on the MMIU meng2024mmiu benchmark, we assess performance across semantic relationships and two-dimensional spatial relationships, as shown in Table 7 and Table 8, respectively. We compare our method with several closed-source models: GPT4o gpt4o , open-source models: Mantis jiang2024mantis , Llava-interleave li2024llavainterleave , and base model Qwen2.5-VL-7B bai2025qwen2 . We achieve the best performance among open-source models across all types.
Ablation Setting Spontaneous Grounding Referential Grounding AVG
CS RS RL Static Robust Common OT MV Region Refer GG Reason Co-Re
Qwen2.5-VL-7B - - - 28.03 5.32 21.81 15.23 5.56 0.54 13.13 32.26 3.96 2.56 14.05
@@ 4.3 Ablation Studies
Effectiveness of Cold Start on GPRO
As discussed in Section 3.2, we directly apply GRPO-based reinforcement learning following the DeepSeek-R1-Zero framework guo2025deepseek . However, this attempt yields only minor performance gains, primarily due to the base model’s constraints in multi-image contexts and complex queries. These limitations underscore the critical importance of implementing a CoT-based cold-start strategy to establish a strong initialization for subsequent reward-based RL learning.
@@
Another key component of our approach is the data filtering via rejection sampling. To assess its impact, we conduct an ablation study by training the stage-1 CoT-SFT model directly with the GRPO algorithm, removing the rejection sampling process. As illustrated in the fourth row of Table 9, our full model—with rejection sampling enabled—achieves a improvement of 3.27%. It shows the importance of data quality in reinforcement learning training and demonstrates the effectiveness of our rejection sampling procedure.

### v2

@@ 4 Experiments
Table 3: In-domain vs Out-of-domain comparison on MIG-Bench li2025migician . The in-domain subsets include static difference, common object recognition, region localization, and referential grounding, while the out-of-domain subsets comprise robust difference, object tracking, multi-view grounding, group grounding, and correspondence. The best result is indicated in bold.
Model Comparison Reasoning Referring $\rm{AP_{50}}$
70B-Scale MLLMs
InternVL2-76B 42.66 35.91 52.78 40.01
InternVL3-78B 45.54 43.11 46.23 43.96
Qwen2.5-VL-72B 71.79 57.95 71.57 64.74
7B-Scale MLLMs
InternVL2-8B 33.90 24.89 34.21 29.28
InternVL3-8B 57.19 37.60 59.05 47.35
Qwen2.5-VL-7B 29.61 20.55 24.23 24.37
+ SFT 60.37 45.02 48.42 52.12
+ CoT-SFT 64.66 45.43 60.89 55.10
Ours 68.93 49.37 59.49 58.49
Table 4: The Comparison on multi-image grounding benchmark MC-Bench xu2024mc . Our method achieve the best performance among 7B-Scale MLLMs, marked in bold.
Model Multi-image Grounding Single image Grounding AVG
ReVOS LISA(val) LISA(test) Ref-L4(val) Ref-L4(test)
Qwen2.5-VL-72B 48.78 53.57 46.85 74.85 76.71 60.15
Qwen2.5-VL-7B 30.07 56.63 53.54 58.57 62.76 52.31
+ SFT 46.33 55.10 54.18 68.92 69.14 58.73
Ours 49.16 62.76 56.11 74.78 75.74 63.71
Table 5: Zero-shot evaluation on several reasoning grounding benchmarks: ReVOS yan2024visa , LISA-Grounding lai2024lisa , and Ref-L4 chen2025revisiting . The best and second-best results are marked in bold and underline, respectively.
GPT4o (Closed-source) 49.17 79.55 31.33 55.33 59.40 72.59 75.00 60.34
Qwen2.5-VL-7B 67.50 45.45 27.33 63.33 51.88 82.96 68.02 58.07
+ SFT 63.33 32.58 26.67 72.67 48.12 90.37 62.21 56.56
Ours 70.00 47.73 32.00 74.67 49.62 84.44 69.77 61.18
Method OP PV WS OPV Vec AVG
@@
Table 7: Object detection evaluation on ODINW li2022grounded dataset. We individually employ five non-overlapping subsets: OP (Oxford Pets), PV (Pascal VOC), WS (Wildfire Smoke), OPV (Open Poetry Vision), and Vec (Vector). The best result is indicated in bold.
4.1 Implementation Details
Benchmarks
We evaluate our method on the multi-image grounding benchmark MIG-Bench li2025migician , MC-Bench xu2024mc , object detection dataset ODINW li2022grounded , single-image grounding datasets: LISA-Grounding lai2024lisa , Ref-L4 chen2025revisiting , and the video grounding dataset ReVOS Grounding yan2024visa . MIG-Bench is a manually curated benchmark designed to evaluate the multi-image grounding ability of MLLMs, which comprises 4.3k instances and covers 10 distinct tasks. MC-Bench features 2K manually annotated samples, containing three distinct styles of text prompts. To align with our experimental setup, we select samples containing one instance per image for evaluation. ODINW comprises 35 distinct real-world settings featuring rare object categories, assessing the model’s object perception and inference ability in practical scenarios. To further evaluate the generalization ability of our model, we individually employ five non-overlapping subsets from ODINW: Oxford Pets, Pascal VOC, Wildfire Smoke, Open Poetry Vision, and Vector. To align with our setting, we select only those samples that contain a single bounding box per image, resulting in a total of 6,314 samples. Ref-L4 is a large-scale (45k annotations) referring expression benchmark with diverse object categories, instance scales, and long textual descriptions, making it ideal for assessing the reasoning capabilities of MLLMs. For ReVOS, we uniformly sample six frames and task the model with grounding one frame, deriving ground-truth bounding boxes from segmentation masks originally annotated for reasoning tasks.
Training Configurations
@@ 4.1 Implementation Details
We select several open-source MLLMs for comparison, such as Qwen2-VL wang2024qwen2 , Qwen2.5-VL bai2025qwen2 , InternVL2 chen2024internvl , InternVL3 zhu2025internvl3 , LLaVA-OneVision li2024llava , MiniCPM2.6 yao2024minicpm , and mPLUG-Owl3 ye2024mplug . To further verify the effectiveness of the RL training, we compare our method with the SFT baseline, which is trained on the same dataset as our stage-1 CoT-SFT with a learning rate of 2e-6 and an accumulated batch size of 64.
@@
Model High-level Sub. High-level Obj. Low-level Two-dimensional Spatial Relationship Average
ISTE HE PT RPM JPS ICSC IQASC ITRSC SOT VGR STS VCor SC FC ER CR MIC VQA FD Sem Spa
GPT4o(Closed-source) 46.5 67.8 93.5 61.3 84.0 72.6 41.5 42.6 79.2 88.8 12.5 23.5 58.5 88.0 17.5 35.0 29.5 55.0 56.0 67.8 41.7
Mantis 31.8 61.5 71.0 65.2 65.5 30.9 28.1 24.3 59.8 57.0 13.5 27.5 48.0 66.0 13.0 23.5 23.0 76.0 50.0 49.5 37.8
Llava-interleave 24.8 29.5 57.0 49.8 54.5 25.1 26.4 23.2 48.8 26.3 15.0 17.5 47.5 42.0 14.0 22.5 21.5 79.0 60.5 36.5 35.5
Qwen2.5-VL-7B 31.0 57.0 80.0 68.0 74.0 44.6 33.6 25.1 82.8 72.0 13.0 26.5 55.5 75.5 7.0 29.0 23.5 58.5 55.5 56.8 37.8
+ SFT 32.5 69.2 87.5 67.0 80.0 40.3 28.6 26.1 81.2 65.1 12.5 30.5 54.0 73.0 19.5 33.5 29.5 43.0 52.5 57.7 38.6
Ours 39.0 69.5 87.0 66.8 72.0 51.1 30.4 27.0 82.2 69.3 15.0 24.5 68.0 83.5 19.0 32.5 22.0 50.0 49.5 59.4 40.4
@@
Table 8: Performance on Multi-image Understanding Benchmark MMIU meng2024mmiu . High-level Obj. and High-level Sub. are represented as subtypes of high-level semantic relationship (objective) and high-level semantic relationship (subjective). The term "Low-level" denotes the subtype of low-level semantic relations. Sem represents the average score across High-level Object, High-level Subject, and Low-level subtype, while Spa denotes the average score for the Two-dimensional Spatial Relationship subtype. Some results are borrowed from the original paper. The best and second-best results among open-source models are marked in bold and underline, respectively.
Ablation Setting Referential Grounding Spontaneous Grounding AVG
CS RS RL OT MV Region Refer GG Reason Co-Re Static Robust Common
Qwen2.5-VL-7B - - - 15.23 5.56 4.07 13.13 32.26 3.96 2.56 28.03 5.32 21.81 13.19
[…]
@@
Ablation Referential Spontaneous AVG
w/o KL 54.20 57.95 55.33
w. KL, $\beta$=1e-1 51.59 57.15 53.26
w. KL, $\beta$=1e-2 54.96 57.89 54.95
w. KL, $\boldsymbol{\beta}$=1e-3 (Ours) 58.16 59.21 58.47
w. KL, $\beta$=1e-4 56.45 58.86 57.17
w. KL, $\beta$=1e-5 52.29 56.98 53.70
Table 10: Ablations for the weight of KL divergence.
4.2 Comparison Results
Multi-image Grounding Evaluation
As shown in Table 2, we present the comparison results in the multi-image grounding benchmark MIG-Bench li2025migician . Our method achieves state-of-the-art performance overall. Specifically, it obtains the best results across all seven subsets of referential grounding. Compared to the base model (Qwen2.5-VL-7B), our approach achieves a 45.64% improvement, and it outperforms the second-best model (Qwen2.5-VL-72B) by 12.60%, while using significantly fewer parameters. We further evaluate our method on MC-Bench xu2024mc , as shown in Table 4. Our model outperforms all 7B-scale MLLMs, the SFT baseline, and the stage-1 CoT-SFT model. Remarkably, it also surpasses several 70B-scale models, including InternVL2-76B and InternVL3-78B, highlighting its strong reasoning and comprehension capabilities.
Moreover, we conduct a comprehensive comparison of various post-training strategies on both in-domain and out-of-domain tasks, shown in Table 3. The results reveal that conventional SFT training significantly improves in-domain performance but provides limited benefits for out-of-domain generalization. CoT-SFT training strategy enhances out-of-domain generalization to a certain extent but slightly compromises in-domain accuracy. Conversely, our full model achieves superior performance, obtaining the highest scores in both in-domain and out-of-domain settings, with a notable 53.34% on out-of-domain tasks. The results demonstrate the strong generalization capability of our approach across both conventional and novel task domains.
Zero-shot Evaluation
To further highlight the robust generalization capability of our method, we conduct zero-shot evaluation on multiple datasets spanning different tasks: the single-image grounding dataset LISA lai2024lisa , Ref-L4 chen2025revisiting , and the video grounding dataset ReVOS yan2024visa . As illustrated in Table 5, our method consistently outperforms all baselines across multiple benchmarks. Although the SFT training strategy achieves slight overall improvements, it exhibits a severe performance drop on the LISA benchmark. It sugguests that the SFT training method lacks strong generalization, relying on memorizing answers instead of learning underlying patterns. On the other hand, our RL training strategy, which is initialized with a small amount of data, achieves a substantial improvement of 4.98% over the SFT baseline and even surpasses the performance of Qwen2.5-VL-72B by 3.56%.
Object Detection Evaluation
The evaluation results on the object detection dataset ODINW are presented in Table 7. Our full model achieves the highest average performance (54.15%) across all five datasets, surpassing the base model Qwen2.5-VL-7B by a significant improvement of 7.24%. Notably, it also outperforms both the SFT baseline and the largest scale base model, Qwen2.5-VL-72B. It demonstrates the robustness and strong generalization capability of our approach in handling out-of-domain object detection tasks.
Multi-image Understanding Evaluation
For comprehensive evaluation on the MMIU meng2024mmiu benchmark, we assess performance across semantic relationships and two-dimensional spatial relationships, as shown in Table 8. We compare our method with several closed-source models: GPT4o gpt4o , open-source models: Mantis jiang2024mantis , Llava-interleave li2024llavainterleave , and base model Qwen2.5-VL-7B bai2025qwen2 . Although our method was not trained on relevant tasks, it achieves the best performance among open-source models across all types.
@@ 4.3 Ablation Studies
Effectiveness of Cold Start
As discussed in Section 3.2, we directly apply GRPO-based RL training following the DeepSeek-R1-Zero framework guo2025deepseek , as shown in the second row of Table 9. However, this attempt yields only minor performance gains (+3.74% over the base model), primarily due to the base model’s constraints in handling multi-image contexts and complex queries. These limitations highlight the critical importance of implementing a CoT cold-start strategy to establish a strong initialization for subsequent reward-based RL training. As presented in the third and fourth rows of Table 9, we employ a CoT cold-start initialization using both full-parameter SFT and LoRA-based SFT, respectively. The results indicate that the cold-start strategy substantially improves the model’s performance. Ultimately, we adopt the LoRA-based cold start approach due to its slightly better performance (+0.59% compared to full-parameter SFT).
@@
Another key component of our approach is data filtering via rejection sampling. To assess its impact, we conduct an ablation study by training the stage-1 CoT-SFT model directly with the GRPO algorithm, removing the rejection sampling process. As illustrated in the fourth row of Table 9, our full model (with rejection sampling enabled) achieves an improvement of 3.27%. It shows the importance of data quality in RL training and demonstrates the effectiveness of our rejection sampling procedure.
Ablations for the weight of KL divergence
Additionally, we conduct an ablation study on the impact of the KL divergence weight during stage-2 RL training. As illustrated in Table 10, removing the KL divergence (w/o KL) results in decreased performance (-3.14%) compared to our full model. Furthermore, we vary the $\beta$ hyperparameter across a range of values [1e-1, 1e-2, 1e-3, 1e-4, 1e-5]. The results indicate that excessively large KL regularization over-constrains the model, while extremely small values provide insufficient constraints. In contrast, an appropriately scaled regularization term, like our 1e-3, offers a balanced trade-off between imposing constraints and encouraging model exploration.
Figure 3: Qualitative comparison. While the Qwen2.5-VL-7B base model and its SFT variant mistakenly grounded the wall and the toilet, respectively, our model correctly understood the instruction and grounded the right object: the shower room.
4.4 Qualitative Visualization
We present a representative case study in Figure 3 to qualitatively evaluate the reasoning and grounding capabilities of our proposed method. The figure demonstrates that while both the Qwen2.5-VL-7B base model and its SFT method fail to accurately ground the target object, our reasoning model successfully accomplishes this task through a coherent two-step reasoning process: (1) identifying the basketball player in Image-1 who requires shower facilities, and (2) precisely localizing the shower room as the appropriate target. In contrast, the base model incorrectly selects the wall, and the SFT method erroneously chooses the toilet. These comparative results provide clear evidence of our model’s enhanced cross-image comprehension and superior instruction-following capabilities.